# Energy landscapes of small SK spin glasses

### v1

@@ Abstract
We study the ±$\pm J$ SK model for small $N$’s up to $N=9$. We sort the $2^{N(N-1)/2}$ possible realizations of the coupling matrix into equivalence classes according to the gauge symmetry and permutation symmetry and determine the energy spectra for each of these classes. We also study the energy landscape in these small systems and find that the elements of the hierarchic organization of ground states start to appear in some samples already for $N$’s as small as 6.
@@ 1 Introduction
The Sherrington-Kirkpatrick model [1], the mean field model of spin glasses, has been the subject of countless studies over the past nearly 50 years. Its solution inspired tremendous progress in a large number of disciplines [2] and earned the 2021 Nobel prize to Giorgio Parisi. It may seem that nothing has been left to learn about the SK model by now. This is not quite true. The overwhelming majority of works about the SK model, including its solution by Parisi [3, 4] and the rigorous work by Guerra [5], Talagrand [6] and others, were all concerned with the thermodynamic limit, where the number $N$ of spins goes to infinity. In this asymptotic limit important simplifications take place, e.g. thermodynamic quantities self-average that is become independent of the realization of disorder. This allows one to approach the problem by averaging over the randomness [7]. Even numerical works on spin glasses, which are necessarily done on finite samples, are performed with quenched averaging and with the ultimate goal of extrapolating to $N\to\infty$ in mind.
@@
However, in a large number of problems where the spin glass metaphor is relevant, the system size is definitely not macroscopic: there are no $10^{24}$ competing agents in an economy, or species fighting for survival in an ecosystem. In several contexts, it may be perfectly justified to study finite, or even small systems. (Think of the example of a family with positive or negative emotional links between the family members who, in addition, are strongly politically committed, left or right, not necessarily overlapping with their emotional relationships. This is evidently a variant of the case of sorting the employees of a firm in two groups in the introduction of Mézard, Parisi and Virasoro’s famous book [8], but $N$ is naturally small in our example.) In such systems self-averaging cannot not hold and quenched averaging cannot be justified; instead, the individual samples have to be studied one by one.
@@
When dealing with a random optimization problem, mathematics is often seeks guarantee against the worst case, while statistical physics is focused on the typical, i.e. the average behavior. We argue that in some real life applications it may be relevant to study the concrete, particular case in question. Of course, the combinatorial explosion of the number of different cases sets a limit to the sizes one can reach in such an endeavor, which makes it important to exploit the symmetries and to classify the different cases into invariance classes, hopefully more manageable in number. This classification is one of the main goals of this paper. In this task we make use of some elementary concepts and results borrowed from the theory of signed graphs [9, 10].
@@
We have also determined the energies and their degree of degeneracy for each of these invariance classes, as well as the corresponding spin configurations. These data allow us to construct the energy landscape at low energies, and also the local energy minima if they appear at higher energies. Because of the huge number of cases, these results are necessarily partial, but illustrate the point sufficiently.
@@
It is often stated that ordering or a phase transition can only take place in an infinitely large system. In practice, already fairly small systems anticipate some characteristic features of the ordered phase in an elementary form if the temperature is low enough. The special sort of order discovered by Parisi [3, 4] in the SK model in the large $N$ limit and on average is complex to an earlier never seen degree. It is interesting to understand how such a complex structure begins to build up in a finite system. Given the messy structure of a disordered system, it is not surprising that, as pointed out by [11], one has to go to very large sizes to observe the emergence of ultrametric geometry based on the metric of overlaps. However, the fragmentation of phase space and the elements of the hierarchic organization of ground states that we regard as the harbingers of the future ultrametric order begin to show up already at $N=6$. We display some examples of this phenomenon in this paper.
@@
A related question is the concentration of measure [6], the fundamental reason behind self-averaging. We find that some invariance classes appear with much larger multiplicity than others when the system size is larger than $N=6$. For example, at $N=9$ the number of independent classes is 1625; out of these 17 appear with multiplicities well beyond a million. Nevertheless, we cannot speak of the dominance of these large-multiplicity classes, because their total multiplicity is still small compared with the total multiplicity $2684354562^{(N-1)(N-2)/2}=2^{28}=$ 268435456 of all the classes - self-averaging is still very very far away.
@@
The plan of the paper is as follows. In Sec.2 we set up the problem and fix notation. In Sec. 3 we make a few remarks and derive some sum rules that are valid for every $N$. In Sec. 4 we present the classification of the configurations and the energy spectra in each class up to $N=6$. Because the number of cases grows extremely fast, the results for $N=7,8$, and 9 cannot be fitted into the paper, so we have set up a site to store this material [12].
@@
In Sec. 5 we exhibit several examples of the energy surface and show how the degenerate ground states divide the phase space into basins of attraction (very small, at these low $N$’s), and how they start to organize into families, the beginning of the evolution toward the ultrametric structure for large $N$’s. The organization of ground states is not the only feature that reflects the lack of convexity of the energy surface: we also find several local minima, saddle points and extended isolated clusters for higher energies. A striking example is a torus consisting of 60 alternating first and second excited states and surrounded by states of higher energy in one of the $N=8$ classes. At the end of Sec. 5, we also display the distribution of ground state energies for $N=6,7,8$, and 9.
@@ 2 Setting up the problem
We study the ±$\pm J$ SK model, the simplest spin glass model placed on the complete graph. The spins will be standard Ising spins, $s_{i}=\pm 1$, $i=1,2,\ldots N$, while the couplings $J$ will also be considered simple binary variables, $J_{i,j}=\pm 1$, $i,j=1,2,\ldots,N$ for $i\neq j$ and $J_{i,i}=0$.
@@
For a given $N$ the number of couplings on the complete graph is $N(N-1)/2$, so the number of different realizations of signed samples is $2^{N(N-1)/2}$. For a given configuration we will have $2^{N}$ spin states ($N$-vectors with components ±$\pm 1$). These numbers grow extremely fast with $N$. In order to investigate the energy landscape, we have to store not only the energies but also the corresponding spin configurations. The volume of data quickly exceeds what can be surveyed, therefore we have to restrict this study to really small systems up to $N=9$. Some of the results we find can, however, be generalized to any $N$.
@@
The couplings $J_{i,j}=J_{j,i}$ and $J_{i,i}=0$ will be defined to take on the values $+1$ and $-1$ in every possible way defining all together $2^{N(N-1)/2}$ different configurations. As we do not plan to play statistical mechanics or consider large systems here, we do not need to use the usual normalization ensuring the extensivity of the energy. This has the advantage of dealing with integers throughout.
@@
$H$ possesses some obvious symmetries that help reducing the number of different cases to be considered. The first such symmetry is the reflection of the spins $s_{i}$ to $-s_{i}$. The Hamiltonian is obviously invariant w.r.t. this transformation.
A reflection of the couplings $J_{i,j}$ to $-J_{i,j}$ results in the reflection of the whole spectrum $\epsilon_{i}$ to $-\epsilon_{i}$ with the degeneracy $\mu_{i}$ remaining the same.
@@
The interaction matrix $J_{i,j}$ as defined above is the adjacency matrix of a signed complete graph. (We are dealing with complete graphs all through this paper.) Signed graphs were introduced in a social psychology context by Harary [9], but the same concepts were discussed earlier under a different terminology in Dénes König’s magisterial book [13].
@@
Finally, the permutation of the vertices collects a number of different $J_{i,j}$’s into equivalence classes in which the energy spectra are identical. The number of different coupling configurations belonging to such a class will be called the multiplicity of the given class. The determination of these multiplicities can be a not entirely trivial combinatorial problem already for moderate $N$’s.
We have also determined the number of frustrated 3 -, 4 -, …$N$-cycles for all the classes up to $N=9$. Remarkably, we found that among the nearly 2000 graphs surveyed there was only a single pair at $N=8$ that had the same numbers of all frustrated $k$-cycles, $k=3,\ldots,N$, but belonged to different invariance classes. Otherwise the number of frustrated cycles was a useful tool to sort the equivalent graphs into invariance classes. The number of cycles will be displayed in the following up to $N=6$ in the main text and to $N=9$ on the web page [12].
@@
Moreover, we have determined the energy spectra and the corresponding spin configurations in all these cases, which are of no concern for workers in the field of signed graphs. We found that, at least up to the size $N=9$, the energy spectra are perfectly capable of distinguishing the different invariance classes, even in the case of the two exceptional $N=8$ graphs where the number of cycles fails to resolve them.
An old paper having some overlap with our results is due to Caliri et al. [16], whose main interest is in the thermodynamics of the system and the zeroes of the partition function on the complex plane. In their Table 1 they list the weights of configurations for $N=5$. Our results for $N=5$ are in agreement with their classification and with their weights that correspond to what we call the multiplicities of these configurations.
3 Some results for general $N$
We can illustrate some results valid for general $N$ on the example of the simple case of $N=4$ where there are only three independent equivalence classes. The corresponding interaction graphs are shown in Fig. 1. For higher $N$’s a complete graph can be very crowded, so we decided to adopt the convention that only the negative edges will be shown, the positive ones left understood. So the first graph in Fig. 1 is in fact the complete graph with all positive edges (the ”ferromagnet”), and it also represents all 8 of its gauge equivalent versions. Similarly, the second and third graph in Fig. 1 are representatives of their 8-member gauge classes. In particular, the third one is gauge equivalent to the purely antiferromagnetic graph, with all its edges negative. Note that we display the representatives of these three classes with the minimal number of negative couplings, as promised. So the minimal signature in these three classes is zero, one, and two, respectively.
@@ 3 Some results for general $N$
The total number of cases for $N=4$ is $642^{N(N-1)/2}=2^{6}=$ 64. As we consider only one representative of each gauge class, we divide this by $82^{(N-1)}=$ 8. So we are left with 8 configurations to distribute among the permutation classes. It is easy to see that we have one purely ferromagnetic and one purely antiferromagnetic configuration, plus six permutation equivalent configurations, corresponding to the number of ways a negative bond can be placed on an N=4 complete graph (that is on the edges of a tetrahedron).
@@
The figures in the first column label the energy levels in this small system, the numbering going from 0 to 6. For general $N$ the possible energy values range from $-\frac{1}{2}N(N-1)$, the ground state energy of the purely ferromagnetic model, to $\frac{1}{2}N(N-1)$, the maximal energy of the pure antiferromagnet. Since the energy steps are 2 units, this makes $\frac{1}{2}N(N-1)+1$ possible energy levels. Note that not each of these energy levels are necessarily occupied in a given class.
The second column shows the energy spectrum in Class 1 with the degeneracy of the energy shown in parentheses. So the ground state energy in Class 1 is -6, twofold degenerate, the energy of the first excited state is zero with a degeneracy of 8, while the highest energy is 2, six times degenerate. In the third column the same data are listed for Class 2, in the fourth the same for Class 3.
For general $N$ the ferromagnetic spectrum goes from $-\frac{1}{2}N(N-1)$ to $\frac{1}{2}N$ for $N$ even, and to $\frac{1}{2}(N-1)$ for $N$ odd. The antiferromagnetic spectrum can be obtained from this by reflection w.r.t. zero. The spectrum of the other classes falls between these two extremes.
@@ 4 The energy spectra
Apart from the smallest $N$’s, making sure that we generate all the configurations and count their multiplicities correctly is not completely trivial. It is expedient to generate them by a systematic method. The one we used, applying a code kindly provided by Matteo Marsili, is as follows: First, we fix the gauge by setting the first line of the interaction matrix as $0,1,1,\dots,1$. This reduces the number of configurations by a factor of $2^{N-1}$, as mentioned before. The remaining matrix elements above the diagonal are laid out as a string of length $(N-1)(N-2)/2$ of ±$\pm{1}$’s. This string is then regarded as a binary number by identifying the matrix element $+1$ with, say, 00 and the $-1$ with 1. Thus simply counting to $2^{(N-1)(N-2)/2}$ guarantees that we generate all the different configurations. Several of these will be equivalent (will have the same spectrum) because of the permutation symmetry, and we have to collect them into equivalence classes. The number of configurations in such a class is the multiplicity of the class. In the following, we will explicitly deal with only one member of the class, the other members being accounted for by the multiplicity.
Unfortunately, generating the configurations as suggested above results in a somewhat random order of their appearance. The first configuration will always be the purely ferromagnetic $J$ (all elements +1), and the last one, the pure antiferromagnetic configuration with all its elements equal to $-1$ (or one of its gauge transforms with less negative matrix elements), but apart from these two extremes the others come out in a more or less random fashion. It is desirable to arrange the different configurations in some logical order. The first idea that may come to mind is to order the configurations according to the number of negative bonds. However, because of the gauge symmetry the number of negative bonds do not provide a consistent classification: graphs containing a different number of negative bonds often belong to the same equivalence class. (We have seen in Fig. 1 that for e.g. $N=4$ a graph with merely two negative edges can stand for one with all six edges being negative.) We found that the lexicographic order according to the number of negative triangles, quadrangles, pentagons, etc. in the complete graph of the interaction matrix provides a possible ordering principle together with an insight into the geometry of interactions.
@@
4.1 Invariance classes and energy spectra for $N$=5

### v3

@@ Abstract
We study the $\pm J$ SK model for small $N$’s up to $N=9$. We sort the $2^{N(N-1)/2}$ possible realizations of the coupling matrix into equivalence classes according to the gauge symmetry and the permutation symmetry and determine the energy spectra for each of these classes. We also study the energy landscape in these small systems and find that the elements of the hierarchical organization of ground states start to appear in some samples already for $N$’s as small as 6.
@@ 1 Introduction
The Sherrington-Kirkpatrick model [1], the mean field model of spin glasses, has been the subject of countless studies over the past nearly 50 years. Its solution inspired tremendous progress in a large number of disciplines [2] and earned the 2021 Nobel prize to Giorgio Parisi. It may seem that nothing has been left to learn about the SK model by now. This is not quite true. The overwhelming majority of works on the SK model, including its solution by Parisi [3, 4] and the rigorous work by Guerra [5], Talagrand [6] and others, were all concerned with the thermodynamic limit, where the number $N$ of spins goes to infinity. In this asymptotic limit important simplifications take place, e.g. thermodynamic quantities self-average, that is become independent of the realization of disorder. This allows one to approach the problem by averaging over the randomness [7]. Even numerical work on spin glasses, which is necessarily done on finite samples, is performed with quenched averaging and with the ultimate goal of extrapolating to $N\to\infty$ in mind.
@@
However, in a large number of problems where the spin glass metaphor is relevant, the system size is definitely not macroscopic: there are no $10^{24}$ competing agents in an economy, or species fighting for survival in an ecosystem. In several contexts, it may be perfectly justified to study finite, or even small systems. (Think of the example of a family with positive or negative emotional links between the family members who, in addition, are strongly politically committed, left or right, not necessarily overlapping with their emotional relationships. This is evidently a variant of the case of sorting the employees of a firm in two groups in the introduction of Mézard, Parisi and Virasoro’s famous book [8], but $N$ is naturally small in our example.) In such systems self-averaging cannot hold and quenched averaging cannot be justified; instead, the individual samples have to be studied one by one.
@@
When dealing with a random optimization problem, mathematics often seeks guarantee against the worst case, while statistical physics is focused on the typical, i.e. the average behavior. We argue that in some real life applications it may be relevant to study the concrete, particular case in question. Of course, the combinatorial explosion of the number of different cases sets a limit to the sizes one can reach in such an endeavor, which makes it important to exploit the symmetries and classify the different cases into classes having identical energy spectra, hopefully more manageable in number. We call these classes isospectral in the following. This classification is one of the main goals of this paper. In this task, we make use of some elementary concepts and results borrowed from the theory of signed graphs [9, 10].
@@
We have also determined the energies and their degree of degeneracy for each of these isospectral classes, as well as the corresponding spin configurations. These data allow us to construct the energy landscape at low energies, and also the local energy minima if they appear at higher energies. Because of the huge number of cases, these results are necessarily partial, but illustrate the point sufficiently.
@@
It is often stated that ordering or a phase transition can only take place in an infinitely large system. In practice, already fairly small systems anticipate some characteristic features of the ordered phase in an elementary form if the temperature is low enough. The special sort of order discovered by Parisi [3, 4] in the SK model in the large $N$ limit and on average is complex to an earlier never seen degree. It is interesting to understand how such a complex structure begins to build up in a finite system. Given the messy structure of a disordered system, it is not surprising that, as pointed out by [11], one has to go to very large sizes to observe the emergence of ultrametric geometry based on the metric of overlaps. However, the fragmentation of phase space and the elements of the hierarchical organization of ground states that we regard as the harbingers of the future ultrametric order start to appear already at $N=6$. In this paper, we display some examples of this phenomenon.
@@
A related question is the concentration of measure [6], the fundamental reason behind self-averaging. We find that some invariance classes appear with much higher multiplicity than others when the system size is larger than $N=6$. For example, at $N=9$ the number of isospectral classes is 1625; out of these 17 appear with multiplicities well beyond a million. Nevertheless, we cannot speak of the dominance of these large-multiplicity classes, because their total multiplicity is still small compared with the total multiplicity $2^{(N-1)(N-2)/2}=2^{28}=268435456$ of all the classes - self-averaging is still very very far away.
@@
The plan of the paper is as follows. In Sec. 2 we set up the problem and fix notation. In Sec. 3 we make a few remarks and derive some sum rules that are valid for every $N$. In Sec. 4 we present the classification of the configurations and the energy spectra in each class up to $N=6$. Because the number of cases grows extremely fast, the results for $N=7,8$, and 9 cannot fit into the paper, so we have set up a site to store this material [12].
@@
In Sec. 5 we exhibit several examples of the energy surface and show how the degenerate ground states divide the phase space into basins of attraction (very small, at these low $N$’s), and how they start to organize into families, the beginning of the evolution toward the ultrametric structure for large $N$’s. The organization of ground states is not the only feature that reflects the lack of convexity of the energy surface: we also find several local minima, saddle points and extended isolated clusters at higher energies. A striking example is a torus consisting of 60 alternating first and second excited states and surrounded by states of higher energy in one of the $N=8$ classes.
@@ 2 Setting up the problem
We study the $\pm J$ SK model, the simplest spin glass model placed on the complete graph. The spins will be standard Ising spins, $s_{i}=\pm 1$, $i=1,2,\ldots N$, while the couplings $J$ will also be considered simple binary variables, $J_{i,j}=\pm 1$, $i,j=1,2,\ldots,N$ for $i\neq j$ and $J_{i,i}=0$.
@@
For a given $N$ the number of couplings on the complete graph is N(N1)/2N(N-1)/2, so the number of different realizations of signed samples is $2^{N(N-1)/2}$. For a given configuration, we will have $2^{N}$ spin states ($N$-vectors with components $\pm 1$). These numbers grow extremely fast with $N$. In order to investigate the energy landscape, we have to store not only the energies but also the corresponding spin configurations. The volume of data quickly exceeds what can be surveyed, therefore we have to restrict this study to really small systems up to $N=9$. However, some of the results we find can be generalized to any $N$.
@@
The couplings $J_{i,j}=J_{j,i}$ and $J_{i,i}=0$ will be defined to take on the values +1+1 and 1-1 in every possible way, defining the $2^{N(N-1)/2}$ different configurations mentioned above. As we do not plan to play statistical mechanics or consider large systems here, we do not need to use the usual normalization ensuring the extensivity of the energy. This has the advantage of dealing with integers throughout.
@@
$H$ possesses some obvious symmetries that help to reduce the number of different cases to be considered. The first such symmetry is the reflection of the spins $s_{i}$ to $s_{i}$. The Hamiltonian is obviously invariant w.r.t. this transformation.
A reflection of the couplings $J_{i,j}$ to $J_{i,j}$ results in the reflection of the whole spectrum $\epsilon_{i}$ to $\epsilon_{i}$ with the degeneracy $\mu_{i}$ remaining the same.
@@
The interaction matrix $J_{i,j}$ as defined above is the adjacency matrix of a signed complete graph. (We are dealing with complete graphs all through this paper.) Signed graphs were introduced in a social psychology context by Harary [9], but the same concepts were discussed earlier under a different terminology in Dénes König’s book [13].
@@
Finally, the permutation of the vertices collects a number of different $J_{i,j}$’s into equivalence classes. Different gauge- and/or permutation-equivalent coupling configurations have the same energy spectra, but the reverse is not necessarily true. When we speak about equivalence classes in the following we always mean isospectral classes, i.e. sets of configurations with the same energy spectra. The number of different coupling configurations belonging to such a class will be called the multiplicity of the given class. The determination of these multiplicities can be a not entirely trivial combinatorial problem already for moderate $N$’s.
We have also determined the number of frustrated 3 -, 4 -, …$N$-cycles for all classes up to $N=9$. Remarkably, we found that among the nearly 2000 graph classes surveyed there was only a single pair at $N=8$ that had the same number of all frustrated $k$-cycles, $k=3,\ldots,N$, but belonged to different invariance classes. Otherwise the number of frustrated cycles was a useful tool to sort the equivalent graphs into invariance classes. The number of cycles will be displayed in the following up to $N=6$ in the main text and for $N=7,8,9$ on the web page [12].
@@
Moreover, we have determined the energy spectra and the corresponding spin configurations in all these cases, which are of no concern for workers in the field of signed graphs.
An old paper having some overlap with our results is due to Caliri et al. [16], whose main interest is in the thermodynamics of the system and the zeros of the partition function on the complex plane. In their Table 1 they list the weights of configurations for $N=5$. Our results for $N=5$ are in agreement with their classification and with their weights that correspond to what we call the multiplicities of these configurations.
Boettcher and Kott [17] consider all the instances of the $\pm J$ SK model and also go up to $N=9$, as we do, but their focus is on the number of frustrated bonds and the distribution of ground state energies, different from ours.
3 Some results for general NN
We can illustrate some results valid for general $N$ on the example of the simple case of $N=4$ where there are only three independent equivalence classes. The corresponding interaction graphs are shown in Fig. 1. For higher $N$’s a complete graph can be very crowded, so we decided to adopt the convention that only the negative edges will be shown, the positive ones left understood. So the first graph in Fig. 1 is in fact the complete graph with all positive edges (the ”ferromagnet”), and it also represents all 8 of its gauge equivalent versions. Similarly, the second and third graph in Fig. 1 are representatives of their 8-member gauge classes. In particular, the third one is gauge equivalent to the purely antiferromagnetic graph, with all its edges negative. Note that we display the representatives of these three classes with the minimal number of negative couplings, as promised. So the minimal signature in these three classes has zero, one, resp. two negative edges.
@@ 3 Some results for general NN
The total number of cases for $N=4$ is $2^{N(N-1)/2}=2^{6}=64$. As we consider only one representative of each gauge class, we divide this by $2^{(N-1)}=8$. So we are left with 8 configurations to distribute among the permutation classes. It is easy to see that we have one purely ferromagnetic and one purely antiferromagnetic configuration, plus six permutation equivalent configurations, corresponding to the number of ways a negative bond can be placed on an $N=4$ complete graph (that is on the edges of a tetrahedron).
@@
The figures in the first column label the energy levels in this small system, the numbering going from 0 to 6. For general $N$ the possible energy values range from $\frac{1}{2}N(N-1)$, the ground state energy of the purely ferromagnetic model, to $\frac{1}{2}N(N-1)$, the maximal energy of the pure antiferromagnet. Since the energy steps are 2 units, this makes $\frac{1}{2}N(N-1)+1$ possible energy levels. Note that not each of these energy levels are necessarily occupied in a given class.
The second column shows the energy spectrum in Class 1 with the degeneracy of the energy shown in parentheses. So the ground state energy in Class 1 is $\penalty 10000\ 6$, two-fold degenerate, the energy of the first excited state is zero with a degeneracy of 8, while the highest energy is 2, six times degenerate. In the third column the same data are listed for Class 2, in the fourth the same for Class 3.
For general $N$ the ferromagnetic spectrum goes from $\frac{1}{2}N(N-1)$ to $\frac{1}{2}N$ for $N$ even, and to $\frac{1}{2}(N-1)$ for $N$ odd. The antiferromagnetic spectrum can be obtained from this by reflection with respect to zero. The spectrum of the other classes falls between these two extremes.
@@ 4 The energy spectra
Apart from the smallest $N$’s, making sure that we generate all the configurations and count their multiplicities correctly is not completely trivial. It is expedient to generate them by a systematic method. The one we used, applying a code kindly provided by Matteo Marsili, is as follows: First, we fix the gauge by setting the first line of the interaction matrix as $0,1,1,\dots,1$. This reduces the number of configurations by a factor of $2^{N-1}$, as mentioned earlier. The upper triangular part of the remaining submatrix is then produced as we count up to $2^{(N-1)(N-2)/2}$ considering the binary representation of the number (counter) as the row-major order of the triangular matrix, associating a positive bond with, say 0 and the negative one with 1. This procedure generates all the different configurations. Several of these will be equivalent (will have the same spectrum) because of the permutation symmetry, and we have to collect them into equivalence classes. The number of configurations in such a class is the multiplicity of the class. In the following, we will explicitly deal with only one member of the class, the other members being accounted for by the multiplicity.
Unfortunately, generating the configurations as suggested above results in a somewhat random order of their appearance. The first configuration will always be the purely ferromagnetic $J$ (all elements +1), and the last one, the pure antiferromagnetic configuration with all its elements equal to 1-1 (or one of its gauge transforms with less negative matrix elements), but apart from these two extremes the others come out in a more or less random fashion. It is desirable to arrange the different configurations in some logical order. The first idea that may come to mind is to order the configurations according to the number of negative bonds. However, because of the gauge symmetry the number of negative bonds do not provide a consistent classification: graphs containing a different number of negative bonds often belong to the same equivalence class. (For example, we have seen in Fig. 1 that for $N=4$ a graph with merely two negative edges can stand for one with all six edges being negative.) We found that the lexicographic order according to the number of negative triangles, quadrangles, pentagons, etc. in the complete graph of the interaction matrix provides a possible ordering principle together with an insight into the geometry of interactions.
@@
4.1 Invariance classes and energy spectra for NN=5